OpenStudy (anonymous):

If 30 percent of 40 percent of a positive number is equal to 20 percent of w percent of the same number, what is the value of w?

OpenStudy (anonymous):

SAY WHAAAAAT? .3*.4*x = (w/100)*x .12*x=w/100*x 12*x=w*x 12=w I GUESS TWELVE AND DOUBLE YOU ARE GAY.